# A Multi-way Parallel Named Entity Annotated Corpus for English, Tamil and Sinhala

### v1

@@ Abstract
journal: Engineering Applications of Artificial Intelligence
@@ 1 Introduction
Named Entity Recognition (NER) is the process of identifying Named Entities (NEs) in natural language text. An NE can be a word or a phrase, and the detected entities are categorized into predetermined categories such as person, location and organization. As an example, consider the sentence: John works in Facebook at Los Angeles. This sentence contains three NEs: John (person), Facebook (organization) and Los Angeles (location). NER either acts as an intermediate step for, or helps to improve many high-level Natural Language Processing (NLP) tasks such as question answering [Lamurias & Couto, 2019], Neural Machine Translation (NMT) [Hu et al., 2022], Information Retrieval [Guo et al., 2009] and automatic text summarization [Khademi & Fakhredanesh, 2020]. There have been recent developments with respect to NE tag sets [Li et al., 2020], tagging schemes [Ringland et al., 2019a], as well as NER algorithms [Li et al., 2020]. Most of the NER algorithms that produced promising results are supervised, meaning that they are trained with NE annotated datasets. However, many low-resource languages have little to no annotated data [Joshi et al., 2020].
@@
Sinhala and Tamil are examples of low-resource languages [Joshi et al., 2020, Ranathunga & de Silva, 2022] with limited NE annotated data [De Silva, 2019].  Manamini et al. [2016]’s is the only publicly available Sinhala NE dataset that has been manually annotated. FIRE corpus [^1] http://fire.irsi.res.in/fire/2023/home is the only publicly available manually annotated NE dataset for Tamil. WikiANN [Pan et al., 2017] and LORELEI[Tracey & Strassel, 2020] are multilingual (but not multi-way pararell) NE annotated datasets that contain both Sinhala and Tamil. However, WikiANN has been automtically annotated using entity linking and LORELEI is hidden behind a paywall.
@@
In this paper, we present a multi-way parallel English-Sinhala-Tamil NE annotated corpus that consists of 3835 sentences per language. This corpus is annotated using the CONLL03 tag set [Sang & De Meulder, 2003], which has four tags: persons (PER), locations (LOC), organizations (ORG) and miscellaneous (MISC). The corpus was annotated using Beginning-Inside-Outside (BIO) format. We provide a comprehensive analysis of this dataset and the data creation process. This dataset is publicly released [^2] https://github.com/suralk/multiNER.
@@
The benefit of having such multi-way parallel datasets is that they can serve as good test beds to evaluate language-specific NER capabilities of pre-trained multilingual Language Models (mLMs) such as mBERT [Devlin et al., 2019] and XLM-R [Conneau et al., 2019], which form the basis of modern-day NLP systems. In other words, if the mLM is fine-tuned with data from individual languages, this results in language-specific NER models, which can be probed to identify how their performance varies across languages. On the other hand, fine-tuning the mLM with all the language data of the multi-way parallel corpus results in a single NER model that caters for all the languages included in the corpus. How the performance of these models varies depending on the linguistic properties of individual languages is useful in building optimal language-specific NER systems.
@@
Despite the NER models built on mLMs outperforming the more traditional Deep Learning (DL) models such as Recurrent Neural Networks (specifically BiLSTM-CRF (Bi-Directional Long Short Term Memory with a CRF layer) [Li et al., 2020, Yadav et al., 2018]), these newer techniques have not been employed for Sinhala or Tamil NER.
@@ 2.1 NE Tag Sets
One prominent tag set that has been widely used in NER is the CONLL03 tag set [Sang & De Meulder, 2003]. This tag set has only 4 tags - Person, Location, Organization and Miscellaneous. The Co-reference and Entity Type tag set [Weischedel & Brunstein, 2005] has 12 NE types (7 numerical types, 10 nominal entity types and temporal types). WNUT2017 [Derczynski et al., 2017] corpus has 6 entity types. ACE 2005/2008 [Song et al., 2015] has identified 7 NE types, as people, organizations, locations, facilities, Geo political entities, weapons, vehicles and events. ACE was defined for the tourism domain, thus can be considered as a domain-specific tag set. While having more NE types helps in extracting more information, training Machine Learning (ML) models for the task becomes very data intensive, which in turn makes manual data annotation costly. As shown in Azeez & Ranathunga [2020a], if a sufficient annotated corpus is not created, having a fine-grained tag set is not useful, as many NE tags will not have sufficient data points to train the ML model.
@@ 2.2 Annotation Schemes
Flat NE annotation schemes have been most commonly used in previous research. Alshammari & Alanazi [2021] present a comprehensive list of such annotation schemes. Some popular NE annotation schemes are IO (tags each token either as an inside tag (I) or an outside tag (O), where NEs are marked as I), IOB (which is also known as BIO - tags each token either as beginning (B) of a known NE, inside (I) it, or outside (O) of any known NE) and IOE (similar to IOB, but tags the end of an NE (E) instead of its beginning).
The main drawback of these flat annotation schemes is that they are unable to capture nested NEs. As an alternative, markup NE annotation schemes have been proposed[Mitchell et al., 2005, Ringland et al., 2019b, Marcus et al., 2011]. While more comprehensive annotation schemes allow the extraction of more fine-grained information, this requires more input from humans who annotate the data and ML models need more data to learn. Therefore, still the BIO tagging scheme, which was used in CONLL 2003 shared task [Sang & De Meulder, 2003] is being commonly used.
@@ 2.3 NER techniques
Recently, pLMs such as BERT [Devlin et al., 2019], RoBERTa  [Liu et al., 2019] and their multilingual variants such as mBERT and XLM-R are widely used for the NER task. In fact, NER was one of the evaluation tasks used for the BERT and XLM-R papers. Subsequent research has introduced various improvements on the vanilla fine-tuning of pLMs for NER [Souza et al., 2019, Fetahu et al., 2022, Huang et al., 2022, Fu et al., 2022]. This line of research has outperformed the traditional DL techniques such as RNNs by a significant margin. Therefore here we discuss only NER techniques based on pLMs. For a comprehensive survey on the previous DL techniques for NER, we refer the reader to Yadav & Bethard [2019].
Some research combined NE corpora from multiple languages (not multi-way parallel) to build a single multilingual NER model on top of a pre-trained multilingual language model [Kulkarni et al., 2023, Shaffer, 2021], which has shown to perform better than language-specific models. This is due to the cross-lingual transfer of knowledge across languages in the model. Thus such multilingual models are a very promising solution for low-resource languages. If there is no NE annotated data for a target language, one solution is to synthetically generate NE data using an NER model of a high-resource language [Li et al., 2021, Yang et al., 2022]. Another popular solution is Transfer Learning, where an NER model fine-tuned on an pLM is used to infer NE information on a target language [Adelani et al., 2021].
All Sinhala NER research except [Rijhwani et al., 2020] is based on traditional Machine Learning (ML) models such as SVM and CRF [Dahanayaka & Weerasinghe, 2014, Manamini et al., 2016, Azeez & Ranathunga, 2020b, Senevirathne et al., 2015, Wijesinghe & Tissera, 2022]Rijhwani et al. [2020] implemented an LSTM based NER model for Sinhala. Similarly for Tamil, except Anbukkarasi et al. [2022], Hariharan et al. [2019], all the other research is based on traditional ML techniques [Murugathas & Thayasivam, 2022, Srinivasagan et al., 2014, Srinivasan & Subalalitha, 2019, Abinaya et al., 2015, Vijayakrishna & Sobha, 2008, Antony & Mahalakshmi, 2014, Abinaya et al., 2014, Gayen & Sarkar, 2014, Murugathas & Thayasivam, 2022, Theivendiram et al., 2018, Mahalakshmi et al., 2016, Malarkodi & Devi, 2020, Malarkodi et al., 2012, Ram et al., 2010].
@@ 2.4 NE Annotated Corpora for Sinhala and Tamil
Manamini et al. [2016] produced an NE annotated dataset [^3] http://bit.ly/2XrwCoK for Sinhala. However, it has not used a standard tagging schema such as BIO. Other Sinhala NER research has not publicly released their datasets. The FIRE corpus is the most commonly used dataset for Tamil NER. Other research [Antony & Mahalakshmi, 2014] that created NER datasets has not published their data.
WikiAnn corpus [Pan et al., 2017] is one of the efforts to build an NE annotated dataset for low resource languages. It was created by transferring annotations from English to other 282 languages through cross-lingual links in knowledge bases. Thus, the WikiAnn corpus is said to have “silver-standard” labels. Although it has both Sinhala and Tamil data, the amount is rather small, specially for Sinhala. Moreover, WikiAnn corpus has been annotated with only three major entities: Person, Organization and Location. LORELEI[Tracey & Strassel, 2020] is another multilingual NE annotated corpus that includes both Sinhala and Tamil. Unlike WikiAnn, LORELEI is a manually annotated corpus. However, it is hidden behind a paywall.
Although Sinhala or Tamil are not included, two other notable multilingual corpora are MultiCoNER [Malmasi et al., 2022] and MasakhaNER [Adelani et al., 2021]. MultiCoNER was built using a technique similar to WikiANN. However data for low-resource languages has been created by Machine Translation. MasakhaNER corpus contains manually annotated data for low-resource African languages. None of these are multi-way parallel.
@@ 3.1 Sinhala
Sinhala is an Indo-Aryan language spoken by about 16 million people, primarily in the island nation Sri Lanka. Sinhala has its own alphabet and script. In Joshi et al. [2020]’s language categorisation, Sinhala is categorized into class 0, meaning that it is a language that has been largely ignored with respect to language technologies and has exceptionally limited resources.
@@ 3.2 Tamil
Tamil is a Dravidian language spoken by about 78 million people, primarily in the Indian state of Tamil Nadu. Tamil is also spoken by a significant population mainly in Sri Lanka, Singapore and Malaysia. Tamil is written in the Tamil script, which is an Abugida (syllabic) script. In Joshi et al. [2020]’s language categorisation, Tamil is categorised into class 3, meaning that it has small amounts of labelled datasets, but with insufficient labelled data collection efforts.
@@ 3.3 Raw Data
This corpus contains duplicate sentences, unwanted long lists (with 200+ tokens per list) such as table of contents and meaningless sentences. Thus, we manually filtered 3835 Sinhala sentences [^4] We started with Sinhala sentences because Fernando et al. [2020]’s corpus was compiled by taking Sinhala as the source.. During the filtering process, we removed meaningless sentences, duplicate sentences, undesirable lists and captions/ headers of figures/tables from the dataset. For each Sinhala sentence in this filtered corpus, we extracted the corresponding Tamil and English sentences from Fernando et al. [2020]’s raw parallel corpus.
@@ 4 Methodology
As mentioned in Section 2.3, the DL techniques used so far for Sinhala and Tamil NER are based on RNN models. Therefore we used one of these models as the baseline. To be specific, we used a BiLSTM CRF model, as it has been the state-of-the-art architecture before the introduction of pre-trained pLMs [Yadav & Bethard, 2019].
@@ 4.1 Bi-LSTM CRF method
We used the BiLSTM-CRF model used by Yadav et al. [2018], which has shown to be the best recurrent model for NER. In addition to character and word embeddings, this model uses prefix-suffix embeddings as the input. The model architecture is shown in Fig. 2. In our implementation, we use FastText to obtain word embeddings. We experimented with both prefix and suffix features, however only affix features resulted in improved results, hence only this result is reported. We used three techniques to generate character embeddings:
@@
    Convolutional Neural Networks (CNNs) [Ma & Hovy, 2016]
@@
    LSTM for character embedding + Affix features[Yadav et al., 2018]
Figure 2: Architecture of the Bi-LSTM CRF network with affix features [Yadav et al., 2018]
@@ 4.2 Fine tuning pre-trained pLMs
Language family-specific pLMs: We experimented with IndicBERT [Kakwani et al., 2020], which has been specifically trained for 12 Indian languages, primarily belonging to Indo-Aryan and Dravidian language families. Tamil is included in indicBERT. Despite being an Indo-Aryan language, Sinhala is not included in this model.
Language-specific pLMs: SinBERT [Dhananjaya et al., 2022] was used as the language-specific pLM for Sinhala. SinBERT was built following RoBERTa. The dataset used to train SinBERT comprises of 15.7 million Sinhala sentences. Note that there is no Tamil-specific pLM.
@@ 6 Evaluation
Table 5 contains the final results. Table 6 shows the tag-wise distribution of results for the best performing model.
@@
First and foremost, we note that Bi-LSTM CRF results with LSTM-generated input embeddings and affix features performed the best among the Bi-LSTM models. However, these results are far below that of pLMs, when the considered language is included in the pre-trained pLM. Note that we did not generate BiLSTM CRF results for English, because the pre-trained models reported the highest results for English.
@@
However, compared to the BiLSTM CRF results, pLMs that have not been pre-trained for a particular language show inferior results for that considered language (e.g. mBERT results for Sinhala). On the positive side, SinBERT, which was trained with just 15.7 million Sinhala sentences, significantly outperforms the corresponding BiLSTM CRF results. This is a promising sign for other low-resource languages as well. Given that many languages have monolingual data, a language-specific pLM can be easily trained and fine-tuned for downstream tasks such as NER.
@@
Similar to Conneau et al. [2019]’s observations, mBERT performs slightly worse than XLM-R for both English and Tamil. On the other hand, IndicBERT falls below the two mLMs. This drop is significant for Tamil. This is very surprising given that IndicBERT was specifically trained on Indic languages, including Tamil and some other languages from the Dravidian family. However, we also note that in Kakwani et al. [2020]’s results IndicBERT fall behind XLM-R and/or mBERT in several NLP tasks. Another interesting fact is SinBERT result for Sinhala, which falls behind XLM-R. A similar observation was made by Dhananjaya et al. [2022] for sentiment analysis. We believe the superior performance of XLM-R is due to its cross-lingual transfer abilities [Asai et al., 2024]. This confirms that mLMs can be a better alternative to language-specific pLMs, specifically when the amount of language-specific data is limited to train pLMs.
@@
There is a noticeable variance across the XLM-R results for the three languages. Given that this is a multi-way parallel dataset (i.e. the same dataset got annotated across the three languages), this variance can be explained with two factors: representation of individual languages in the mLM, and complexity of individual languages. For example, English, which has the highest representation in XLM-R (i.e. English language data has been used much more than other languages during XLM-R model pre-training) has shown the highest performance. However, despite Tamil having more representation in XLM-R than Sinhala, XLM-R result for Tamil is much lower than Sinhala. This could be due to the language complexity (Tamil is more agglutinative than Sinhala). Another reason could be Dravidian languages (where Tamil belongs to) being under-represented in XLM-R compared to the Indo-Aryan languages (where Sinhala belongs to) [Ranathunga & de Silva, 2022].
@@
Finally, we note that multilingually fine-tuned XLM-R (mXLM-R) performs the best for both Sinhala and Tamil. However, there is a slight drop in the result for English. We believe this is due to the phenomenon known as negative interference, which states that the model performance for high-resource languages drops in multilingual settings [Wang et al., 2020]. Nevertheless, since this single model can handle all three languages, this multilingual model is the best pick for our task.

### v2

@@ Abstract
journal: Natural Language Processing
@@ 1 Introduction
Named Entity Recognition (NER) is the process of identifying Named Entities (NEs) in natural language text. An NE can be a word or a phrase, and the detected entities are categorized into predetermined categories such as person, location and organization. As an example, consider the sentence: John works in Facebook at Los Angeles. This sentence contains three NEs: John (person), Facebook (organization) and Los Angeles (location). NER either acts as an intermediate step for, or helps to improve many high-level Natural Language Processing (NLP) tasks such as question answering [1], Neural Machine Translation (NMT) [2], Information Retrieval [3] and automatic text summarization [4]. There have been recent developments with respect to NE tag sets [5], tagging schemes [6], as well as NER algorithms [5]. Most of the NER algorithms that produced promising results are supervised, meaning that they are trained with NE annotated datasets. However, many low-resource languages have little to no annotated data [7].
@@
Sinhala and Tamil are examples of low-resource languages [7, 8] with limited NE annotated data [9]. Manamini et al.’s [10] is the only publicly available Sinhala NE dataset that has been manually annotated. FIRE corpus [^1] http://fire.irsi.res.in/fire/2023/home is the only publicly available manually annotated NE dataset for Tamil. WikiANN [11] and LORELEI[12] are multilingual (but not multi-way pararell) NE annotated datasets that contain both Sinhala and Tamil. However, WikiANN has been automtically annotated using entity linking and LORELEI is hidden behind a paywall.
@@
In this paper, we present a multi-way parallel English-Sinhala-Tamil NE annotated corpus that consists of 3835 sentences per language. This corpus is annotated using the CONLL03 tag set [13], which has four tags: persons (PER), locations (LOC), organizations (ORG) and miscellaneous (MISC). The corpus was annotated using Beginning-Inside-Outside (BIO) format. We provide a comprehensive analysis of this dataset and the data creation process. This dataset is publicly released [^2] https://github.com/suralk/multiNER.
@@
The benefit of having such multi-way parallel datasets is that they can serve as good test beds to evaluate language-specific NER capabilities of pre-trained multilingual Language Models (mLMs) such as mBERT [14] and XLM-R [15], which form the basis of modern-day NLP systems. In other words, if the mLM is fine-tuned with data from individual languages, this results in language-specific NER models, which can be probed to identify how their performance varies across languages. On the other hand, fine-tuning the mLM with all the language data of the multi-way parallel corpus results in a single NER model that caters for all the languages included in the corpus. How the performance of these models varies depending on the linguistic properties of individual languages is useful in building optimal language-specific NER systems.
@@
Despite the NER models built on mLMs outperforming the more traditional Deep Learning (DL) models such as Recurrent Neural Networks (specifically BiLSTM-CRF (Bi-Directional Long Short Term Memory with a CRF layer) [5, 16]), these newer techniques have not been employed for Sinhala or Tamil NER.
@@ 2.1 NE Tag Sets
One prominent tag set that has been widely used in NER is the CONLL03 tag set [13]. This tag set has only 4 tags - Person, Location, Organization and Miscellaneous. The Co-reference and Entity Type tag set [17] has 12 NE types (7 numerical types, 10 nominal entity types and temporal types). WNUT2017 [18] corpus has 6 entity types. ACE 2005/2008 [19] has identified 7 NE types, as people, organizations, locations, facilities, Geo political entities, weapons, vehicles and events. ACE was defined for the tourism domain, thus can be considered as a domain-specific tag set. While having more NE types helps in extracting more information, training Machine Learning (ML) models for the task becomes very data intensive, which in turn makes manual data annotation costly. As shown by Azeez and Ranathunga [20], if a sufficient annotated corpus is not created, having a fine-grained tag set is not useful, as many NE tags will not have sufficient data points to train the ML model.
@@ 2.2 Annotation Schemes
Flat NE annotation schemes have been most commonly used in previous research. Alshammari et al. [21] present a comprehensive list of such annotation schemes. Some popular NE annotation schemes are IO (tags each token either as an inside tag (I) or an outside tag (O), where NEs are marked as I), IOB (which is also known as BIO - tags each token either as beginning (B) of a known NE, inside (I) it, or outside (O) of any known NE) and IOE (similar to IOB, but tags the end of an NE (E) instead of its beginning).
The main drawback of these flat annotation schemes is that they are unable to capture nested NEs. As an alternative, markup NE annotation schemes have been proposed [22, 23, 24]. While more comprehensive annotation schemes allow the extraction of more fine-grained information, this requires more input from humans who annotate the data and ML models need more data to learn. Therefore, still the BIO tagging scheme, which was used in the CONLL 2003 shared task [13] is being commonly used.
@@ 2.3 NER techniques
Recently, pLMs such as BERT [14], RoBERTa  [25] and their multilingual variants such as mBERT and XLM-R are widely used for the NER task. In fact, NER was one of the evaluation tasks used for the BERT and XLM-R papers. Subsequent research has introduced various improvements on the vanilla fine-tuning of pLMs for NER [26, 27, 28, 29]. This line of research has outperformed the traditional DL techniques such as RNNs by a significant margin. Therefore here we discuss only NER techniques based on pLMs. For a comprehensive survey on the previous DL techniques for NER, we refer the reader to Yadav et al. [30].
Some research combined NE corpora from multiple languages (not multi-way parallel) to build a single multilingual NER model on top of a pre-trained multilingual language model [31, 32], which has shown to perform better than language-specific models. This is due to the cross-lingual transfer of knowledge across languages in the model. Thus such multilingual models are a very promising solution for low-resource languages. If there is no NE annotated data for a target language, one solution is to synthetically generate NE data using an NER model of a high-resource language [33, 34]. Another popular solution is Transfer Learning, where an NER model fine-tuned on an pLM is used to infer NE information on a target language [35].
All Sinhala NER research except Rijhwani et al. [36] who implemented an LSTM based NER model for Sinhala is based on traditional Machine Learning (ML) models such as SVM and CRF [37, 10, 38, 39, 40]. Similarly for Tamil, except Anbukkarasi et al. [41] and Hariharan et al. [42], all the other research is based on traditional ML techniques [43, 44, 45, 46, 47, 48, 49, 50, 43, 51, 52, 53, 54, 55].
@@ 2.4 NE Annotated Corpora for Sinhala and Tamil
Manamini et al. [10] produced an NE annotated dataset [^3] http://bit.ly/2XrwCoK for Sinhala. However, it has not used a standard tagging schema such as BIO. Other Sinhala NER research has not publicly released their datasets. The FIRE corpus is the most commonly used dataset for Tamil NER. Other research [48] that created NER datasets has not published their data.
WikiAnn corpus [11] is one of the efforts to build an NE annotated dataset for low resource languages. It was created by transferring annotations from English to other 282 languages through cross-lingual links in knowledge bases. Thus, the WikiAnn corpus is said to have “silver-standard” labels. Although it has both Sinhala and Tamil data, the amount is rather small, specially for Sinhala. Moreover, WikiAnn corpus has been annotated with only three major entities: Person, Organization and Location. LORELEI [12] is another multilingual NE annotated corpus that includes both Sinhala and Tamil. Unlike WikiAnn, LORELEI is a manually annotated corpus. However, it is hidden behind a paywall.
Although Sinhala or Tamil are not included, two other notable multilingual corpora are MultiCoNER [56] and MasakhaNER [35]. MultiCoNER was built using a technique similar to WikiANN. However, data for low-resource languages has been created by Machine Translation. MasakhaNER corpus contains manually annotated data for low-resource African languages. None of these are multi-way parallel.
@@ 3.1 Sinhala
Sinhala is an Indo-Aryan language spoken by about 16 million people, primarily in the island nation Sri Lanka. Sinhala has its own alphabet and script. In Ranathunga and de Silva’s [8] language categorisation, Sinhala is categorized into class 2. According to Joshi et al.’s [7] language category definition, this means Sinhala has only a small amount of annotated datasets.
@@ 3.2 Tamil
Tamil is a Dravidian language spoken by about 78 million people, primarily in the Indian state of Tamil Nadu. Tamil is also spoken by a significant population mainly in Sri Lanka, Singapore and Malaysia. Tamil is written in the Tamil script, which is an Abugida (syllabic) script. In Ranathunga and de Silva’s [8] language categorisation, Tamil is categorised into class 3, meaning that it has small amounts of labelled datasets, but with a better web presence than Sinhala.
@@ 3.3 Raw Data
This corpus contains duplicate sentences, unwanted long lists (with 200+ tokens per list) such as table of contents and meaningless sentences. Thus, we manually filtered 3835 Sinhala sentences [^4] We started with Sinhala sentences because Fernando et al.’s [57] corpus was compiled by taking Sinhala as the source.. During the filtering process, we removed meaningless sentences, duplicate sentences, undesirable lists and captions/ headers of figures/tables from the dataset. For each Sinhala sentence in this filtered corpus, we extracted the corresponding Tamil and English sentences from Fernando et al. [57]’s raw parallel corpus.
@@ 4 Methodology
As mentioned in Section 2.3, the DL techniques used so far for Sinhala and Tamil NER are based on RNN models. Therefore we used one of these models as the baseline. To be specific, we used a Bi-LSTM CRF model, as it has been the state-of-the-art architecture before the introduction of pre-trained pLMs [30].
@@ 4.1 Bi-LSTM CRF method
We used the BiLSTM-CRF model used by Yadav et al. [16], which has shown to be the best recurrent model for NER. In addition to character and word embeddings, this model uses prefix-suffix embeddings as the input. The model architecture is shown in Fig. 2. In our implementation, we used FastText to obtain word embeddings. We experimented with both prefix and suffix features, however only prefix features resulted in improved results, hence only this result is reported. We used three techniques to generate character embeddings:
@@
    Convolutional Neural Networks (CNNs) [58]
@@
    LSTM for character embedding + Affix features[16]
Figure 2: Architecture of the Bi-LSTM CRF network with affix features [16]
@@ 4.2 Fine tuning pre-trained pLMs
Language family-specific pLMs: We experimented with IndicBERT [60], which has been specifically trained for 12 Indian languages, primarily belonging to Indo-Aryan and Dravidian language families. Tamil is included in indicBERT. Despite being an Indo-Aryan language, Sinhala is not included in this model.
Language-specific pLMs: SinBERT [61] was used as the language-specific pLM for Sinhala. SinBERT was built following RoBERTa. The dataset used to train SinBERT comprises of 15.7 million Sinhala sentences. Note that there is no Tamil-specific pLM.
@@ 6 Evaluation
Table 5 contains the final results. Table 6 shows the tag-wise distribution of results for the XLM-R model.
@@
First and foremost, we note that Bi-LSTM CRF results with LSTM-generated input embeddings and prefix features performed the best among the Bi-LSTM models. However, these results are far below that of pLMs, when the considered language is included in the pLM. Note that we did not generate Bi-LSTM CRF results for English, because the pLMs reported the highest results for English.
@@
However, compared to the Bi-LSTM CRF results, pLMs that have not been pre-trained for a particular language show inferior results for that considered language (e.g. mBERT results for Sinhala). On the positive side, SinBERT, which was trained with just 15.7 million Sinhala sentences, significantly outperforms the corresponding Bi-LSTM CRF results. This is a promising sign for other low-resource languages as well. Given that many languages have monolingual data, a language-specific pLM can be easily trained and fine-tuned for downstream tasks such as NER.
@@
Similar to Conneau’s [15] observations, mBERT performs slightly worse than XLM-R for both English and Tamil. On the other hand, IndicBERT falls below the two mLMs. This drop is significant for Tamil. This is very surprising given that IndicBERT was specifically trained on Indic languages, including Tamil and some other languages from the Dravidian family. However, we also note that in Kakwani et al.’s [60] results, IndicBERT falls behind XLM-R and/or mBERT in several NLP tasks. Another interesting fact is SinBERT result for Sinhala, which falls behind XLM-R. A similar observation was made by Dhananjaya et al. [61] for sentiment analysis. We believe the superior performance of XLM-R is due to its cross-lingual transfer abilities [62]. This confirms that mLMs can be a better alternative to language-specific pLMs, specifically when the amount of language-specific data is limited to train pLMs.
@@
There is a noticeable variance across the XLM-R results for the three languages. Given that this is a multi-way parallel dataset (i.e. the same dataset got annotated across the three languages), this variance can be explained with two factors: representation of individual languages in the mLM, and complexity of individual languages. For example, English, which has the highest representation in XLM-R (i.e. English language data has been used much more than other languages during XLM-R model pre-training) has shown the highest performance. However, despite Tamil having more representation in XLM-R than Sinhala, XLM-R result for Tamil is much lower than Sinhala. This could be due to the language complexity (Tamil is more agglutinative than Sinhala). Another reason could be Dravidian languages (where Tamil belongs to) being under-represented in XLM-R compared to the Indo-Aryan languages (where Sinhala belongs to) [8].
@@
Finally, we note that multilingually fine-tuned XLM-R (mXLM-R) performs the best for both Sinhala and Tamil. However, there is a slight drop in the result for English. We believe this is due to the phenomenon known as negative interference, which states that the model performance for high-resource languages drops in multilingual settings [63]. Nevertheless, since this single model can handle all three languages, this multilingual model is the best pick for our task.